# Efficient Artifacts Removal for Adaptive Deep Brain Stimulation and a Temporal Event Localization Analysis

### v1

@@ 3.2.2. Prepare data for ANN
Instead of processing all segments in $\operatorname{\boldsymbol{X}}_{m}$ at once, they were divided into smaller groups, each containing 500 segments, and the LFP was cleaned up by eOptShrink. This grouping strategy was implemented to speed up the eOptShrink algorithm. For the sake of self-completeness, eOptShrink algorithm can be found in A. After LFP removal by eOptShrink, each segment underwent wavelet transformation using the Haar wavelet and we obtain the wavelet coefficient matrix:
@@
where $q\in\mathbb{N}$ is the number of wavelet coefficients. The average stimulation amplitude, denoted as $a_{m}:=\frac{1}{n_{m}}\sum_{j=1}^{n_{m}}\max x_{j}$, and the median stimulus artifact, denoted as $s_{m}\in\mathbb{R}^{q}$, were computed as key characteristics of the $m$-th recording for subsequent analysis. Feature selection for applying ANN and KNN was then performed as follows. First, a matrix of all median stimulus artifacts $s_{m}$ was constructed: $S=[s_{1}\ s_{2}\ldots s_{M}]\in\operatorname{\mathbb{R}}^{q\times(M+1)}$. For each $i=1,\ldots,q$, the variance of the $i$th row of $S$ was computed and denoted as $s_{v}(i)$, yielding a variance vector $s_{v}\in\mathbb{R}^{q}$. The entries of $s_{v}$ were then sorted in descending order, and their cumulative sum was computed, resulting in a vector $\tilde{s}_{v}\in\mathbb{R}^{q}$. The index $k$ was defined as the largest integer such that $\tilde{s}_{v}(k)<0.99\tilde{s}_{v}(q)$. The indices of the top $k$ entries in the sorted array were then selected as features for the subsequent ANN and KNN analyses.
3.3. Build decision trees
@@
For the $(M+1)$th aDBS signal, find recordings with stimulation amplitude different from $a_{M+1}$ by less than 10%; that is, all $m$ s.t. $\left|\frac{a_{m}-a_{M+1}}{a_{M+1}}\right|\leq 0.1$ were selected as candidates. Among these candidates, the KNN algorithm was applied to identify $Q$ recordings with closest $s_{m}$ to $s_{M+1}$. If fewer than $Q$ candidates were available, all candidates were included. Collect all stimulus artifact as $\operatorname{\mathbb{G}}_{M+1}=\left[w_{1}\ w_{2}\ \ldots\ w_{n^{\prime}}\right]\in\mathbb{R}^{q\times n^{\prime}}$, where $n^{\prime}$ represents the number of segments in the selected $K$ recordings.
@@
Decision trees were then constructed using $\operatorname{\mathbb{G}}_{M+1}$. To build a tree, two points $w_{i}$, $w_{j}$ were randomly selected to generate a hyperplane, which is determined by the middle point $\frac{w_{i}+w_{j}}{2}$ with the normal direction $\frac{w_{i}-w_{j}}{\|w_{i}-w_{j}\|}$. The data was then divided into two subspaces by the hyperplane. Once one subspace contains fewer than $M^{\prime}$ points, that subspace is not further touched. Otherwise, this process was repeated until all subspaces contained fewer than $M^{\prime}$ elements, and hence a tree structure. The resulting subspaces were termed end nodes of the tree, and the data points they contained were referred to as leaves of the tree. Since the selection of points for hyperplane derivation was random, the results from a single decision tree were unstable and potentially inaccurate. To enhance accuracy, $T$ trees were constructed, forming a decision forest. While a larger $T$ improved accuracy and stability, it also increased computational cost.
@@ 3.4. Derive template and recover LFPs
For each artifact segment, the corresponding wavelet coefficients was obtained. These coefficients were then used to locate the corresponding leaves in the decision forest. Specifically, for each node in a decision tree, the subspace corresponding to the wavelet coefficient set was determined by the hyperplane of that node. The set continued traversing the tree until it reached an end node, where the contained leaves were considered its neighbors. After searching through the entire forest, the ANN algorithm identified at most $M\times N$ neighbors, collectively denoted as $\operatorname{\boldsymbol{Y}}$.
Next, the KNN was applied to $\operatorname{\boldsymbol{Y}}$, selecting $K$ closest neighbors. The final template for the $i$th stimulus artifact, denoted as $\hat{\boldsymbol{s}}^{*}_{i}$, was then computed by taking the entry-wise median of the $K$ chosen neighbors. The optimal value of $K$ was determined using a greedy search algorithm that minimized the artifact residual (AR) index [26] defined as
(3) $\begin{split}\mathrm{AR}_{i}=&\left|\mathrm{log}\left(\frac{1}{2}\left[\frac{\mathrm{med}|\hat{Z}_{i}-\mathrm{med}(\hat{Z}_{i})|}{\mathrm{med}|Z_{i}-\mathrm{med}(Z_{i})|}+\frac{\mathrm{med}|Z_{i}-\mathrm{med}(Z_{i})|}{\mathrm{med}|\hat{Z}_{i}-\mathrm{med}(\hat{Z}_{i})|}\right]\right)\right.\\ &\left.\times\frac{1}{2}\left[\frac{\mathrm{Q95}|\hat{Z}_{i}-\mathrm{med}(\hat{Z}_{i})|}{\mathrm{Q95}|Z_{i}-\mathrm{med}(Z_{i})|}+\frac{\mathrm{Q95}|Z_{i}-\mathrm{med}(Z_{i})|}{\mathrm{Q95}|\hat{Z}_{i}-\mathrm{med}(\hat{Z}_{i})|}\right]\right|,\end{split}$
where $\hat{Z}_{i}$ represents the first $t_{1}$ ms and $Z_{i}$ the last $t_{2}$ ms of the estimated LFP signal $\boldsymbol{x}_{i}-\hat{\boldsymbol{s}}_{i}^{*}$ obtained after template subtraction. The functions $\mathrm{med}$ and $\mathrm{Q95}$ denote the median and 95th percentile, respectively. In this study, $t_{1}=3$ ms and $t_{2}=5.2$ ms. The optimal $K$ was selected from the candidate values $K=10,20,\ldots,100$ to minimize the AR index.
@@
where $g_{1}$ and $g_{2}$ represent the number of overlapping points with the preceding and following segments, respectively. The final artifact-free LFP signal was obtained by subtracting the sequence of artifact templates from the original signal. Additionally, signal trend removal was performed using the smoothing and polynomial fitting method described in Sec. 3.1. Notably, stimulation periods could be identified prior to artifact removal.
@@ 3.5.1. Semi-real aDBS recording generation
With a large $K$, not all selected neighbors closely resembled the target artifacts, but this approach effectively removed the underlying LFP. The overlap-and-add method was then applied to reconstruct the semi-real artifacts, and the previously extracted trends for both stimulation and non-stimulation periods were reintroduced. As a result, the generated signal retained the stimulus artifacts and the trend of the original aDBS signal while minimizing the presence of underlying LFP activity. Finally, the semi-real aDBS signal was obtained by combining the artifact-only signal with a LFP signal recorded in the absence of stimulation.
@@ 3.5.2. Evaluation measures
The effectiveness of artifact removal was quantified using three indices: artifact residual (AR), spectral concentration (SC), and normalized mean square error (NMSE), as described in [24]. The AR for the entire signal was computed by averaging the AR index, $\mathrm{AR}_{i}$, across all segments (i=1,,Ni=1,...,N, where $N$ was the total number of segments). The AR index for each segment, $\mathrm{AR}_{i}$, was defined in Eq. 3, where $Z_{i}$ was formed by concatenating the last $t_{2}$ ms of the (i20)(i-20)th segment to the $(i+20)$th segment.
@@
The NMSE was calculated as
@@
For an aDBS algorithm, two key concerns were: (1) how accurately beta events were detected and (2) how many detected events were incorrect. To carry out the temporal event localization analysis, consider the following metrics. The first one was evaluated using recall, defined as
(7) $\mathrm{Recall}=\frac{\#\mathrm{TP}}{\#\mathrm{TP}+\#\mathrm{FN}}.$
The second concern was assessed using precision, defined as
(8) $\mathrm{Precision}=\frac{\#\mathrm{TP}}{\#\mathrm{TP}+\#\mathrm{FP}}.$
These two indices were further summarized using the F1-score, given by
(9) $\mathrm{F1}=\frac{2\times\mathrm{Recall}\times\mathrm{Precision}}{\mathrm{Recall}+\mathrm{Precision}}.$
Recall quantified the proportion of true events detected, precision measured the proportion of detected events that are true, and the F1-score provided a balanced summary of recall and precision.
@@ 3.5.3. Other methods for comparison
To validate the performance of SMARTA+, comparative analyses were conducted against SMARTA and two artifact removal techniques: template subtraction and blanking. SMARTA was implemented following the methodology described in [24], without any modifications. Template subtraction was implemented based on the causal setup described in [7], where templates are generated in real time using only past data. Specifically, for the $i$th segment, ensuring that the estimation was computed as the average of the (ik)(i-k)th to (i1)(i-1)th segments, ensuring that the estimation relies solely on preceding information. Notably, although [7] employed a hybrid approach combining both front-end and back-end modules, only the back-end component of template subtraction could be implemented in this study due to data constraints.
Blanking was also implemented in two forms: pulse blanking and transient blanking. Front-end blanking, as commonly applied in embedded aDBS devices [32], was realized using the sample-and-interpolate technique [14] to perform pulse blanking. Each stimulation artifact pulse was blanked over a 2.5 ms window, with the removed samples interpolated to maintain continuity. Transient blanking, implemented according to the method proposed in [1], was applied at the level of entire stimulation periods rather than individual artifact pulses. Specifically, a 550 ms window following the onset of each stimulation period was blanked to suppress the broadband noise associated with DC transient artifacts.
@@ 4. Results
The SMARTA+ algorithm was implemented in MATLAB 2019b and executed on a machine with an AMD Ryzen 5 3500X 6-core processor, 16 GB RAM, and Windows 10. In this study we have $M=34$. For SMARTA+, the parameters were set to $Q=5$, $M^{\prime}=500$, $T=50$.
@@ 4.1. Visualization of different algorithms
SMARTA+ was evaluated in both time and frequency domains and compared with SMARTA, template subtraction, and blanking. As shown in Figure 4, although the stimulus artifact signal (Figure 4E) lacks beta activity (13-35 Hz), trend-induced spectral spreading contaminated the beta band. Moreover, transient DC artifacts introduced beta-like activity at stimulation onsets and offsets (Figure 4D), distorting the beta amplitude in the semi-real signal.
@@
In the time domain, artifacts near the stimulation onset differed from those during stable stimulation (Figure 4BC). As a result, SMARTA struggled to find sufficient matching neighbors within the same signal, leading to suboptimal artifact templates. In contrast, due to the help of ANN, SMARTA+ utilized neighbors from other recordings, allowing for more appropriate template generation. While both methods performed similarly during stable stimulation periods (Figure 4C), as shown in Figure 4B, SMARTA+ (red line) outperformed SMARTA (blue line) at the stimulation onsets since SMARTA+ estimated and removed transient DC artifacts within stimulation periods. In the frequency domain (Figure 4E), both SMARTA and SMARTA+ successfully removed the aliasing peak at 42.8 Hz, but SMARTA+ further reduced low-frequency noise compared to SMARTA since the transient DC artifact is removed and stimulus artifact near onset is better recovered. Regarding beta amplitude estimation, SMARTA failed to fully estimate contamination at stimulation onsets and offsets (Figure 4D). In contrast, SMARTA+ achieved better artifact removal at onsets and successfully removed trends after offsets, allowing the base LFP’s beta amplitude to be recovered more accurately.
Figure 4. Validation of SMARTA+ on a semi-real aDBS signal. (A) Waveform of the generated artifact signal (dashed line), semi-real aDBS signal (black line), base LFP (green line), and artifact-free LFPs obtained using SMARTA (blue line) and SMARTA+ (red line). (B) Zoomed-in view of artifacts at the onset of stimulation. (C) Zoomed-in view of artifacts during stable stimulation. (D) Beta amplitude estimated using the aDBS algorithm. (E) Power spectrum derived using Welch’s method.
Figure 5 presents a comparison using a real aDBS signal, demonstrating results consistent with those observed in the semi-real case. SMARTA struggled to fully remove onset artifacts and DC trends in the time domain, while SMARTA+ effectively eliminated them. In the frequency domain, SMARTA+ achieved greater suppression of low-frequency noise. Beta amplitudes were significantly distorted at the stimulation onsets and offsets; SMARTA partially reduced these effects, but SMARTA+ was able to largely remove them.
@@
Figure 5. Validation of SMARTA+ on a real aDBS signal. (A) Waveform of the real aDBS signal (black line), and artifact-free LFPs obtained using SMARTA (blue line) and SMARTA+ (red line). (B) Zoomed-in view of artifacts at the onset of stimulation. (C) Zoomed-in view of artifacts during stable stimulation. (D) Beta amplitude estimated using the aDBS algorithm. (E) Power spectrum derived using Welch’s method.

### v2

@@ 3.2.2. Prepare data for ANN
Instead of processing all segments in $\operatorname{\boldsymbol{X}}_{m}$ at once, they were divided into smaller groups, each containing $2p$ segments, and the LFP was cleaned up by eOptShrink. This grouping strategy was implemented to speed up the eOptShrink algorithm. For the sake of self-completeness, eOptShrink algorithm can be found in A. After LFP removal by eOptShrink, each segment underwent wavelet transformation using the Haar wavelet and we obtained the wavelet coefficient matrix:
@@
where $q\in\mathbb{N}$ is the number of wavelet coefficients. The average stimulation amplitude, denoted as $a_{m}:=\frac{1}{n_{m}}\sum_{j=1}^{n_{m}}\max x_{j}$, and the median stimulus artifact, denoted as $s_{m}\in\mathbb{R}^{q}$, were computed as key characteristics of the $m$-th recording for subsequent analysis. Feature selection for applying ANN and KNN was then performed as follows. First, a matrix of all median stimulus artifacts $s_{m}$ was constructed: $S=[s_{1}\ s_{2}\ldots s_{M}]\in\operatorname{\mathbb{R}}^{q\times M}$. For each $i=1,\ldots,q$, the variance of the $i$th row of $S$ was computed and denoted as $s_{v}(i)$, yielding a variance vector $s_{v}\in\mathbb{R}^{q}$. The entries of $s_{v}$ were then sorted in descending order, and their cumulative sum was computed, resulting in a vector $\tilde{s}_{v}\in\mathbb{R}^{q}$. The index $k$ was defined as the largest integer such that $\tilde{s}_{v}(k)<0.99\tilde{s}_{v}(q)$. The indices of the top $k$ entries in the sorted array were then selected as features for the subsequent ANN and KNN analyses.
3.3. Build decision forest
@@
For the $(M+1)$th aDBS signal, recordings were selected from the other $M$ recordings with stimulation amplitude different from $a_{M+1}$ by less than 10%; that is, all $m$ s.t. $\left|\frac{a_{m}-a_{M+1}}{a_{M+1}}\right|\leq 0.1$ were selected as candidates. Among these candidates, the KNN algorithm was applied to identify $Q$ recordings with closest $s_{m}$ to $s_{M+1}$. If fewer than $Q$ candidates were available, all candidates were included. Collect all stimulus artifacts as $\operatorname{\mathbb{G}}_{M+1}=\left[w_{1}\ w_{2}\ \ldots\ w_{n^{\prime}}\right]\in\mathbb{R}^{q\times n^{\prime}}$, where $n^{\prime}$ represents the number of segments in the selected $K$ recordings.
@@
Decision trees were then constructed using $\operatorname{\mathbb{G}}_{M+1}$. To build a tree, two points $w_{i}$, $w_{j}$ were randomly selected to generate a hyperplane, which is determined by the middle point $\frac{w_{i}+w_{j}}{2}$ with the normal direction $\frac{w_{i}-w_{j}}{\|w_{i}-w_{j}\|}$. The data were then divided into two subspaces by the hyperplane. Once a subspace contained fewer than $M^{\prime}$ points, it was not further partitioned. Otherwise, this process was repeated until all subspaces contained fewer than $M^{\prime}$ elements, thus yielding a tree structure. The resulting subspaces were termed end nodes of the tree, and the data points they contained were referred to as leaves of the tree. Since the selection of points for hyperplane derivation was random, the results from a single decision tree could be unstable and potentially inaccurate. To enhance accuracy, $T$ trees were constructed, forming the initial decision forest. While a larger $T$ improved accuracy and stability, it also increased computational cost.
During the first stimulation period, SMARTA+ employed the initial decision forest. Upon completion of each stimulation period, artifacts extracted from that period were added to the artifact library, after which the decision forest was reconstructed online according to the method outlined previously. This iterative update enabled the forest to gradually adapt to the target signal.
@@ 3.4. Derive template and recover LFPs
For each artifact segment, the corresponding wavelet coefficients were obtained. These coefficients were then used to locate the corresponding leaves in the decision forest. Specifically, for each node in a decision tree, the subspace corresponding to the wavelet coefficient set was determined by the hyperplane of that node. The set continued traversing the tree until it reached an end node, where the contained leaves were considered its neighbors. After searching through the entire forest, the ANN algorithm identified at most $M\times N$ neighbors, collectively denoted as $\operatorname{\boldsymbol{Y}}$.
Next, the KNN was applied to $\operatorname{\boldsymbol{Y}}$, selecting $K$ closest neighbors. The final template for the $i$th stimulus artifact, denoted as $\hat{\boldsymbol{s}}^{*}_{i}$, was then computed by taking the entry-wise median of the $K$ chosen neighbors. The optimal value of $K$ was determined using a greedy search algorithm that minimized the artifact residual (AR) index [malik2017ARindex], defined as
(3) $\begin{split}\mathrm{AR}_{i}=&\left|\mathrm{log}\left(\frac{1}{2}\left[\frac{\mathrm{med}|\hat{Z}_{i}-\mathrm{med}(\hat{Z}_{i})|}{\mathrm{med}|Z_{i}-\mathrm{med}(Z_{i})|}+\frac{\mathrm{med}|Z_{i}-\mathrm{med}(Z_{i})|}{\mathrm{med}|\hat{Z}_{i}-\mathrm{med}(\hat{Z}_{i})|}\right]\right)\right.\\ &\left.\times\frac{1}{2}\left[\frac{\mathrm{max}|\hat{Z}_{i}-\mathrm{med}(\hat{Z}_{i})|}{\mathrm{max}|Z_{i}-\mathrm{med}(Z_{i})|}+\frac{\mathrm{max}|Z_{i}-\mathrm{med}(Z_{i})|}{\mathrm{max}|\hat{Z}_{i}-\mathrm{med}(\hat{Z}_{i})|}\right]\right|,\end{split}$
where $\hat{Z}_{i}$ represents the first $t_{1}$ ms and $Z_{i}$ the last $t_{2}$ ms of the estimated LFP signal $\boldsymbol{x}_{i}-\hat{\boldsymbol{s}}_{i}^{*}$ obtained after template subtraction. The functions $\mathrm{med}$ and $\mathrm{max}$ denote the median and the maximum, respectively. In this study, $t_{1}=3$ ms and $t_{2}=5.2$ ms. The optimal $K$ was selected from the candidate values $K=5,15,\ldots,55$ to minimize the AR index.
@@
where $g_{1}$ and $g_{2}$ represent the number of overlapping points with the preceding and following segments, respectively. The final artifact-free LFP signal was obtained by subtracting the sequence of artifact templates from the original signal. Additionally, signal trend removal was performed using the smoothing and polynomial-fitting method described in Sec. 3.1. Notably, stimulation periods could be identified prior to artifact removal.
@@ 3.5.1. Semi-real aDBS recording generation
With a large $K$, not all selected neighbors closely resembled the target artifacts, but this approach effectively removed the underlying LFP. The overlap-and-add method was then applied to reconstruct the semi-real artifacts, and the previously extracted trends for both stimulation and non-stimulation periods were reintroduced. As a result, the generated signal retained the stimulus artifacts and the trend of the original aDBS signal while minimizing the presence of underlying LFP activity. Finally, the semi-real aDBS signal was obtained by combining the artifact-only signal with an LFP signal recorded in the absence of stimulation.
In clinical adaptive DBS systems, the neural sensing sampling rate is typically around 1 kHz [cummins2021DBSfs]. Accordingly, the semi-real recordings were resampled to 1.1 kHz to evaluate SMARTA+ under a clinic-like acquisition setting.
To further assess SMARTA+’s capability of handling different clinical stimulation parameters, the stimulation frequency was adjusted. Motivated by stimlation parameters used in DBS [cagnan2019DBS180], after constructing the semi-real artifacts as described above, artifact-only signals at a 180 Hz stimulation rate were added.
@@ 3.5.2. Evaluation measures
The effectiveness of artifact removal was quantified using three indices: artifact residual (AR), spectral concentration (SC), and normalized mean square error (NMSE), as described in [liu2024]. The AR for the entire signal was computed by averaging the AR index, $\mathrm{AR}_{i}$, across all segments (i=1,,Ni=1,...,N, where $N$ was the total number of segments). The AR index for each segment, $\mathrm{AR}_{i}$, was defined in Eq. 3, where $Z_{i}$ was formed by concatenating the last $t_{2}$ ms of the (i20)(i-20)th segment to the $(i+20)$th segment.
@@
In the semi-real setup, the NMSE was calculated as
@@
For an aDBS algorithm, two key concerns were: (1) how accurately beta events were detected and (2) how many detected events were incorrect. To carry out the temporal event localization analysis, consider the following metrics. The first one was evaluated using recall, defined as $\mathrm{Recall}=\frac{\#\mathrm{TP}}{\#\mathrm{TP}+\#\mathrm{FN}}$. The second concern was assessed using precision, defined as $\mathrm{Precision}=\frac{\#\mathrm{TP}}{\#\mathrm{TP}+\#\mathrm{FP}}$. These two indices were further summarized using the F1-score, given by $\mathrm{F1}=\frac{2\times\mathrm{Recall}\times\mathrm{Precision}}{\mathrm{Recall}+\mathrm{Precision}}$. Recall quantified the proportion of true events detected, precision measured the proportion of detected events that are true, and the F1-score provided a balanced summary of recall and precision.
@@ 3.5.3. Other methods for comparison
To validate the performance of SMARTA+, comparative analyses were conducted against SMARTA and two artifact removal techniques: template subtraction and blanking. SMARTA was implemented following the methodology described in [liu2024], without any modifications. Template subtraction was implemented based on the causal setup described in [culaclii2016ts], where templates are generated in real time using only past data. Specifically, for the $i$th segment, the template was computed as the average of the (ik)(i-k)th to (i1)(i-1)th segments, ensuring that the estimation relied solely on preceding information. Notably, although [culaclii2016ts] employed a hybrid approach combining both front-end and back-end modules, only the back-end component of template subtraction could be implemented in this study due to data constraints.
Blanking was also implemented in two forms: pulse blanking and transient blanking. Front-end blanking, as commonly applied in embedded aDBS devices [stanslaski2018aDBSdevice], was realized using the sample-and-interpolate technique [heffer2008interpolation] to perform pulse blanking. Each stimulation artifact pulse was blanked over a 2.5 ms window, with the removed samples interpolated to maintain continuity. Transient blanking, implemented according to the method proposed in [anso2022artifactRemoval], was applied at the level of entire stimulation periods rather than individual artifact pulses. Specifically, a 550 ms window following the onset of each stimulation period was blanked to suppress the broadband noise associated with DC transient artifacts.
@@ 4. Results
The SMARTA+ algorithm was implemented in MATLAB 2020a and executed on a machine with an AMD Ryzen 5 3500X 6-core processor, 16 GB RAM, and Windows 10. In this study we have 34 recordings. For SMARTA+, the parameters were set to $Q=3$, $M^{\prime}=500$ and $T=50$. To ensure general applicability, no parameter tuning was performed. Benchmark algorithms were implemented with their recommended parameter settings.
@@ 4.1. Visualization of different algorithms
Building on prior reports that SMARTA outperforms template subtraction and blanking [liu2024], SMARTA+ was examined in both time and frequency domains for enhanced visualization and compared with SMARTA in this subsection. As shown in Figure 4D, trend-induced spectral spreading contaminated the beta band. Moreover, transient DC artifacts introduced beta-like activity at stimulation onsets and offsets (Figure 4C), distorting the beta amplitude, which was derived using the aDBS protocol described in Sec. 2.1, in the semi-real signal.
In the time domain, SMARTA+ attenuated but did not fully eliminate stimulation artifacts in the first stimulation period because the initial library and decision forest came from out-of-signal exemplars (from other recordings), resulting in a template discrepancy. After each stimulation period, artifacts from the just-processed period were appended to the library and the decision forest was retrained, enabling template estimation from in-signal artifacts. This online adaptation made SMARTA+ progressively more stable and effective. As shown in Figure 4A, artifacts persisted in the first period but were effectively removed in the subsequent stimulation periods. A closer examination of the third stimulation period, shown in Fig. 4B, indicates that both SMARTA and SMARTA+ reliably reconstructed the underlying LFP signals, demonstrating stable performance after the initial adaptation phase, which typically requires only a few stimulation periods (approximately three).
@@
In the frequency domain (Figure 4D), both SMARTA and SMARTA+ successfully removed the aliasing peak at 42.8 Hz, but SMARTA+ further reduced low-frequency noise compared to SMARTA since transient DC artifacts were removed and onset-related stimulus artifacts were better recovered. Regarding beta amplitude estimation, SMARTA failed to fully estimate contamination at stimulation onsets and offsets (Figure 4C). In contrast, SMARTA+ achieved better artifact removal at onsets and successfully removed trends after offsets, allowing the base LFP’s beta amplitude to be recovered more accurately.
@@
Figure 4. Validation of SMARTA+ on a semi-real aDBS signal. Periods without stimulation are omitted to improve visualization. (A) Waveform of the semi-real aDBS signal (gray line), base LFP (black line), and artifact-free LFPs obtained using SMARTA (blue line) and SMARTA+ (red line). (B) Zoomed-in view of artifacts at the third stimulation period. (C) Beta amplitude estimated using the aDBS algorithm. (D) Power spectrum derived using Welch’s method.
Figure 5 presents a comparison using a real aDBS signal, demonstrating results consistent with those observed in the semi-real case. SMARTA+ was unable to fully remove artifacts during the first two stimulation periods, after which its performance stabilized and remained effective. In the frequency domain, SMARTA+ achieved greater suppression of low-frequency noise. Beta amplitudes were significantly distorted at the stimulation onsets and offsets; SMARTA partially reduced these effects, whereas SMARTA+ substantially attenuated them.
Figure 5. Validation of SMARTA+ on a real aDBS signal. Periods without stimulation are omitted to improve visualization. (A) Waveform of the real aDBS signal (gray line), and artifact-free LFPs obtained using SMARTA (blue line) and SMARTA+ (red line). (B) Zoomed-in view of artifacts at the third stimulation period. (C) Beta amplitude estimated using the aDBS algorithm. (D) Power spectrum derived using Welch’s method.